Operations Report – January 27th

Operations Report

Crew 220 Operations Report 27JAN2020

Sol: 8

Name of person filing report: Matt Wise

Non-nominal systems: Primary Generator, Water Heater, Hab lower deck fluorescent lights

Generator: Primary is inoperable. Rental was run-the following information is for the rental.

Hours run: 14hrs
From what time last night: 18:00

To what time this morning: 08:00

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): 92% Not yet run

Diesel Reading – Below 1/2 (approx 48%)

Station Propane Reading- 82%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 7

Water Meter (units)- 0148360,6

Water (static tank) (gal) – 201

Static to loft Pump used – Yes

Water in GreenHab (gal): 150.34

Water in Science Dome (gal): 0

Toilet Tank Emptied – Yes

Deimos rover used: Still in the workshop


Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: No


Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used – No

Starting Hours: 127.9

Beginning Charge: 100%

Ending Hours: 128

Ending Charge: 100%

Opportunity Rover used: Still in workshop


Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: No

Starting Hours: 133.9

Beginning Charge: 100%

Ending Hours: 134

Ending Charge: 100%

Notes on rovers: Opportunity and Deimos off-site for maintenance. There were no EVAs today, so the rovers were not used.

ATVs Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? No

# hours the ATV’s were used today: 0

Notes on ATVs: None.

Habcar used and why, where? Yes; Post Office & Supplies

CrewCar used and why, where? No

General Notes and Comments: None.

Summary of the internet: Nominal

Summary of Suit and Radios: 3 radios (black w/ blue tape) malfunctioning; one doesn’t power on, one transmits off/on rapidly when keyed, and one scrambles/screen flashes when keyed. Team conducted basic troubleshooting, power cycle, and reset settings with no effect.
Suits Nominal

Summary of Hab: Main fluorescent lights in Hab lower deck dim and flickering. Center light bar usually lighting correctly, light near door flickers and dim all times, light bar near stairs has no bulbs. Engineer attempted to remove/replace bulbs with no effect. We’re not aware of any replacement bulbs in Hab or RAM. Any further troubleshooting must be left to qualified professional

Summary of Science Dome operations: Nominal

Summary of Ram operations: Nominal

Summary of any observatory issues: Nominal

Summary of Health and Safety issues: MDRS crew all healthy

Questions, concerns, and requests to Mission Support: None. Thanks!

Copyright © The Mars Society. All rights reserved. | Main Site